A car radiator contains 10 liters of a 30% antifreeze solution. How many liters will have to be replaced with pure antifreeze if the resulting solution is to be 50% antifreeze?
Answer by nerdybill(7384)

A car radiator contains 10 liters of a 30% antifreeze solution. How many liters will have to be replaced with pure antifreeze if the resulting solution is to be 50% antifreeze?
.
Let x = amount (liters) replaced with pure antifreeze
then
.30(10-x) + x = .50(10)
3-.30x + x = 5
3+.70x = 5
.70x = 2
x = 2/.70
x = 20/7 liters
or, approximately
x = 2.86 liters